OpenStudy (anonymous):

What does (y4z2)(y3z2) equal? Explain using the expanded form and the multiplication property of exponents

OpenStudy (anonymous):

y^7z^4

OpenStudy (anonymous):

Since exponent values are added when their bases are multiplied together, it becomes (y4z2)(y3z2) = y^(4+3) * z^(2+2) = y^7z^4

OpenStudy (anonymous):

so when you imagine in a^b * a^c = a*a*a... (b times) * a*a*a...(c times) so now the number of "a"s multiplied together is b+c thus a*(b+c) so a*b=a*c=a*(b+c)